# Deformable Cluster Manipulation via Whole-Arm Policy Learning

### v1

@@ VI LIMITATIONS AND FUTURE WORK
A few directions for future work to reduce the sim-to-real gap and improve the success rates are as follows. First, the quality of perception can be enhanced with multiple camera views or with partial-to-complete point cloud models for surface reconstruction. Second, to introduce additional real-world complexities during simulation, integrating with visual tree reconstruction works [44] that identifies the stem structure beneath the leaf clutter, or system identification methods [29] that infers parameters governing branch dynamics, may be necessary. Third, external whole-arm tactile sensors [10] could be leveraged instead of our proprioceptive-only model to enhance the real-world applicability and effectiveness of contact-aware policies. Finally, the generalisability of our approach needs to be tested in out-of-context distributions, for example, with cables, ropes, or sutures.

### v2

@@ VI LIMITATIONS AND FUTURE WORK
A few directions for future work to reduce the sim-to-real gap and improve the success rates are as follows. First, the quality of perception can be enhanced with multiple camera views or with partial-to-complete point cloud models for surface reconstruction. Second, to introduce additional real-world complexities during simulation, integrating with system identification methods [jacob2024learning] that infers parameters governing branch dynamics, may be necessary. Third, external whole-arm tactile sensors [huang20243d] could be leveraged instead of our proprioceptive-only model. Finally, our approach could extend beyond agriculture; for instance, to clear deformable cable bundles obstructing inspection cameras in data-centres, reposition flexible ducting during aircraft maintenance, or manipulate soft tissue clusters under occlusion in robotic surgery using whole-arm contact.